A wave is modeled with the function y(x,t) = (0.46m) cos((0.72rad/m)x − (0.77rad/s)t + π/3). Find the period.




Round your answer to 2 decimal places.



Do not forget the unit

1
Expert's answer
2022-05-13T03:24:43-0400
"y(x,t) = (0.46m) \cos((0.72rad/m)x \\− (0.77rad/s)t + π/3)"

"y(x,t) = y_0 \cos(2\pi/\lambda)x − (2\pi/T)t + \phi_0)"

"2\pi/T=0.77\:\rm rad/s"

"T=2\pi/(0.77\:\rm rad/s)=8.16\: s"